OpenStudy (anonymous):

The following ODE is solved by the following function. How do I determine the values of a, b and c?

OpenStudy (anonymous):

ODE: \[y' + ax^b \sin (x^c) y =0\] Soluioin: \[y = e^{ \cos (x^4)}\]

OpenStudy (anonymous):

Should I do this: Calculate y' from the solution, thus \[y' = e^{ \cos (x^4)} \times - \sin (x^4) \times 4x^3\] Now I plug this into the ODE: \[\color {green} a x^{\color {red} b} \sin (x^{\color {blue} c}) y =\color {green}4x^{\color {red} 3} \sin (x^{\color {blue} 4}) e^{ \cos (x^4)}\] and get that a = 4, b = 3 c = 4?

OpenStudy (luigi0210):

Is this calculus?

OpenStudy (luigi0210):

Well this high school senior just got through calculus one so I won't be much help here. My apologies

OpenStudy (anonymous):

What would y' be? I did it again and got my old answer again

OpenStudy (anonymous):

@Luigi0210 hahahaah:)

OpenStudy (luigi0210):

I will observe and learn

OpenStudy (anonymous):

I did it right?

OpenStudy (luigi0210):

What? oh darn it -_-

OpenStudy (anonymous):

@Loser66 cool! I'm currently busy with Cal3 and some other subjects. next semester is cal4 (i think numberical analysis?) Oh and sorry, but I don't see what I did wrong?

OpenStudy (loser66):

oh, yeah, since you put - out side of sin, it confused me, Oh I am sorry. my mis read, i think you got y' = ... -(minus) .... my bad, I apology.!!

OpenStudy (loser66):

@taljaards in my college, differential equations is a separate course after cal3.

OpenStudy (phi):

yes, I get a=4, b=3, c=4

OpenStudy (anonymous):

@Loser66 haha you confused me too haha, but nvm thanks for the help! Thanks @phi! At ours Cal3 also isn't DE... I have DE, Cal3, linear algebra1 and other subjects..